OpenStudy (anonymous):

(-2)(-3)(0)= 0, right? and -4/0= 0 again?

OpenStudy (anonymous):

oh, and thats -4 over 0

OpenStudy (anonymous):

you can't divide by zero

OpenStudy (anonymous):

the first one is 0 the second is undefined

OpenStudy (freckles):

yes to first no to second

OpenStudy (freckles):

you cannot have division by 0

OpenStudy (anonymous):

any number over 0 is undefined

OpenStudy (anonymous):

so are the two equal, or no?

OpenStudy (unklerhaukus):

multiples of zero are zero but you cannot divide by zero \[{-4 \over 0}\] is not defined

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\mathbb{R}/0 = undefined\] I recommend remembering this as it becomes important in calculus and functions

OpenStudy (anonymous):

\[\mathbb{R}\] being all real numbers

OpenStudy (anonymous):

ok, thanks

Directrix (directrix):

Division is defined in terms of multiplication. Think about 6/ 3 or 6 divided by 3. 6/3 = 2 if 2 * 3 = 6 Apply that idea to -4/0. -4 divided by 0 = 0 if 0 * 0 = -4. It does not. That may help in remembering that division by zero is undefined. Note that 0/4 = 0 because 0 * 4 = 0